OpenStudy (anonymous):

The quantity of carbon-14 in an organic sample decreased by 1% in a period of 83 years. Use this information to estimate the half-life of carbon-14.

OpenStudy (anonymous):

so after 83 years, there is 99 percent left let A be the initial amount .99A = A(1/2)^(t/halflife) ,

OpenStudy (anonymous):

The general halflife equation is y = A (1/2)^(t/H) where A is initial amount and H is halflife after 83 years, there is 99 percent left .99A = A(1/2)^(83/H) .99 = (1/2)^(83/H) ln .99 = ln ((1/2)^(83/H)) ln .99 = 83/H *ln (1/2) , solve for H H = 83 * ln (1/2)/ ln(.99)

OpenStudy (anonymous):

H = 5724.3 years

OpenStudy (anonymous):

Cheers

OpenStudy (anonymous):

did that make sense?